OpenStudy (anonymous):

What is the slope of the line whose equation is –48 = 2x − 8y? as a fraction

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

type and ill give u1uri

OpenStudy (anonymous):

i cant give u one in chat

OpenStudy (uri):

8y=2x 48 y=(2x+48)/8 y=1/4x+6

OpenStudy (uri):

Solve =1/4